if perimeter of a square is 150, find the area
Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

150/4=37.5
37.5^2=1406.25 sq units

Answer by mananth(16946)

Perimeter= 150
square edge be x
all are equal
so 4x=150
x=37.5 the side length
..
Area = L*W
But it is a square
so Area = L*L
Are = 37.5*37.5
Area = 1406.25